Use the accompanying data set on the pulse rates (in beats per minute) of males to complete parts (a) and (b) below. Click thX- Evaluate z = if x= 32.6, u = 49.3, 6 = 9.3, and n=23. a To Z= (Type an integer or decimal rounded to two decimal places asFind the integers that correspond to the given description. fewer than 25 people Select the correct choice below and fill inX- Evaluate za if x = 65, = 74.9, and o = 2.3. z (Type an integer or decimal rounded to two decimal places as needed.)
Answer #1

Solni, @) From the table which We are given n = 42 and Exi=2936 We know that Exi 2936 Mean = 42 where Exi= sum of all pluse rto our answer is (A) O A. The integers up to 25 (d) Evaluate Z= X-M # / if x=65 M=74-9 ord o=2-3 but the value in equation 2